(xz+yz)dz/dy + (xz-yz)dz/dy = x2+y2 find the general solution


1
Expert's answer
2021-02-01T08:14:16-0500

"(xz(y)+yz(y))z'(y)+(xz(y) - yz(y))z'(y)=x^2+y^2"

Ordinary differential equation classification:

first order nonlinear ordinary differential equation

"\frac{\large dz(y)}{\large dy}(xz(y)-yz(y))+\frac{\large dz(y)}{\large dy}(xz(y)+yz(y))=x^2+y^2"

Solve for "\frac{\large dz(y)}{\large dy}"

"\space\space\frac{\large dz(y)}{\large dy} = \frac{\large x^2+y^2}{\large 2xz(y)}"

Multiply both sides by "2xz(y):"

"2x\frac{\large dz(y)}{\large dy}z(y)=x^2+y^2"

Integrate both sides with respect to y:

"\int 2x\frac{\large dz(y)}{\large dy}z(y)dy=\int (x^2+y^2)dy"

Evaluate the integrals:

"xz(y)^2 = \frac{\large y^3}{\large 3}+x^2y+C_1."       where "C_1" is an arbitrary constant.

Solve for z(y):

"Answer: \space z(y) = - \large\frac{\sqrt{y^3+3x^2y+3C_1}}{\sqrt{3x}} \space or \space \large\frac{\sqrt{y^3+3x^2y+3C_1}}{\sqrt{3x}}"